# Task-Adaptive Pretrained Language Models via Clustered-Importance Sampling

### v1

@@ Abstract
Specialist language models (LMs) focus on a specific task or domain on which they often outperform generalist LMs of the same size. However, the specialist data needed to pretrain these models is only available in limited amount for most tasks. In this work, we build specialist models from large generalist training sets instead. We adjust the training distribution of the generalist data with guidance from the limited domain-specific data. We explore several approaches, with clustered importance sampling standing out. This method clusters the generalist dataset and samples from these clusters based on their frequencies in the smaller specialist dataset. It is scalable, suitable for pretraining and continued pretraining, it works well in multi-task settings. Our findings demonstrate improvements across different domains in terms of language modeling perplexity and accuracy on multiple-choice question tasks. We also present ablation studies that examine the impact of dataset sizes, clustering configurations, and model sizes.
@@ 2 Related Work
Generalist vs Specialist LMs Generalist LMs address tasks which they have not been trained for explicitly (Brown et al., 2020), or provide a good initialization for fine-tuning a dedicated model (Devlin et al., 2019). Nowadays generalists compete with dedicated models on many tasks (Jiang et al., 2024; Dubey et al., 2024). Success however comes at a price: a generalist must be much larger than a specialist for the same accuracy. For instance, on English-to-German translation, the 175-B parameter generalist GPT-3 (Brown et al., 2020) is less accurate than a 136m-parameter specialist (Sennrich et al., 2016a). For neural LMs, the parameter count directly impacts training and inference costs.
Specialist large LMs exist in domains where large amounts of specialized texts are available. Corpora with billions of tokens enable pretraining or continued pretraining, a generalist pretraining phase followed by a specialist one (Gururangan et al., 2020; Parmar et al., 2024)). Domains with specialist models include medicine and biology (Lewis et al., 2020; Labrak et al., 2024; Bolton et al., 2024), computer programming and mathematics (Lewkowycz et al., 2022; Rozière et al., 2024; Azerbayev et al., 2024) and finance (Wu et al., 2023; Xie et al., 2023a). When specialist data is available in limited amount, there are methods to train specialist models on generalist data instead.
Task-Adaptive Data-Selection The selection methods over-sample generalist data that aids model generalization in the target domain. For masked LMs, Gururangan et al. (2020) observe that continued pretraining improves the performance on end-tasks when using data with high vocabulary overlap with the targeted task. For machine translation (MT), Aharoni & Goldberg (2020) show that a task-adapted pretraining dataset can be selected from a generalist dataset using the nearest neighbors of a small specialist set. Their nearest neighbor classifier relies on BERT sentence distance (Devlin et al., 2019). Still for MT, other works have used other types of classifiers. In particular, contrasting the scores of two LMs (generalist and specialist) is popular (Moore & Lewis, 2010; Axelrod et al., 2011; Wang et al., 2018; Junczys-Dowmunt, 2018). Other classifiers include logistic regression or fine-tuned BERT (Iter & Grangier, 2021). Outside classification, Xie et al. (2023c) proposed to use importance sampling for continued pretraining. They improve classification tasks by selecting pretraining data with a similar distribution to the targeted domain in terms of hashed-ngrams. Importance sampling is also used in (Grangier et al., 2024b) and we build upon that work which adjusts the frequency of generalist clusters informed by specialist data: we scale the method to millions of clusters, show that it works with larger models, and extend it beyond language modeling tasks.
@@ 3.1 Notations
$D^{\rm g}$ is the training dataset sampled from the generalist distribution ${\cal D}^{\rm g}$. $D^{\rm s}$ is the specialist dataset representative of the final task, sampled from ${\cal D}^{\rm s}\neq{\cal D}^{\rm g}$. The loss of model $\theta$ on $D$ is
@@
where $|D|$ denotes the cardinality of a set $D$ and $|x|$ denotes the length of sequence $x=x^{|x|}_{0}=(x_{1},\ldots,x_{|x|})$. The perplexity of model $\theta$ on the dataset $D$ is ${\cal P}(D;\theta):=\exp({\cal L}(D;\theta)).$
@@ 3.2 Classification
which restrict the generic dataset $D^{\rm g}$ to the examples with an estimated probability to be in-domain above threshold $t$. The threshold is validated as a trade-off between focusing on data close to the domain of interest while keeping $C(D_{g},t)$ large enough to train an LM of the targeted capacity. In our case, we rely on a logistic regression classifier trained over sentence BERT (SBERT) text embeddings (Reimers & Gurevych, 2019), an established classification method (Minaee et al., 2021). The SBERT representation is also commonly used in data selection (Albalak et al., 2024; Xie et al., 2023c; Zhang et al., 2024; Su et al., 2023). This representation is also used in the alternative selection strategies we consider. As an ablation, we also evaluate the impact of the choice of SBERT (Section 5.1).
@@ 3.4 CRISP: ClusteRed Importance Sampling for Pretraining
Figure 1: Clustered-based Importance Sampling with CRISP.
@@
${\cal L}({\cal D}^{\rm s};\theta)=\sum_{c}{\cal L}(c;\theta)P(c|{\cal D}^{\rm s})=\sum_{c}{\cal L}(c;\theta)\frac{P(c|{\cal D}^{\rm s})}{P(c|{\cal D}^{\rm g})}P(c|{\cal D}^{\rm g})=\mathop{\mathbb{E}}_{c\sim(c|{\cal D}^{\rm s})}[w(c){\cal L}(c;\theta)].$
In our experiments, we estimate the terms $w(c),{\cal L}(c;\theta)$ from the finite training sets $D^{\rm s}\sim{\cal D}^{\rm s}$ and $D^{\rm g}\sim{\cal D}^{\rm g}$. We count the number of examples in each cluster to estimate $P(c|{D}^{\rm s}),P(c|{D}^{\rm g})$. The expected loss over a cluster ${\cal L}(c;\theta)$ is estimated as the average loss over the generalist examples in cluster $c$, ${\cal L}({D}^{\rm g}\cap K(c);\theta)$, where $K(c)$ denotes the examples in cluster $c$. This strategy therefore only estimates $P(c|{D}^{\rm s})$ on the small ${D}^{\rm s}$. The term ${\cal L}(c;\theta)$ is estimated over the large set as ${\cal L}({D}^{\rm g}\cap K(c);\theta)$ which has less variance than the estimator ${\cal L}({D}^{\rm s}\cap K(c);\theta)$ over the small ${D}^{\rm s}$.
We train CRISP models with stochastic optimization (Kingma & Ba, 2015, Adam) and propose Algorithm 1. Here, we do not explicitly reweigh the loss. We instead samples cluster from their importance. This avoids frequently visiting clusters with less weights. This strategy has less variance in its gradient estimates, which can help convergence (Seiffert et al., 2008; An et al., 2021). This algorithm is simple and efficient when one groups the generalist examples by cluster prior to training.
@@
8:         $x_{i}\sim{\rm Uniform}({D}^{\rm s}\cap K(c))$triangleright Sample a generalist example in the selected cluster.
@@
10:     $\theta_{t}\leftarrow{\rm AdamUpdate}(\theta_{t-1},\{x_{0},\ldots,x_{B}\})$
@@ 4 Experiments & Results
We perform experiments with transformer LMs (Vaswani et al., 2017). Most of our experiments use models with 1.3B parameters (trained on 120B tokens) and we conduct ablations with 350m and 7B models (resp. trained on 40B, 350B tokens). Our settings for architectures and optimization are borrowed from Brown et al. (2020), see Appendix C.
Our generalist training set is Redpj2 (Together AI Team, 2023). We select this dataset as it contains only web-crawled data without additional interventions to help evaluation tasks (e.g. adding encyclopedias, books or academic articles). Redpj2 contains over 30T tokens with our 32k byte-pair encoding tokenizer (Sennrich et al., 2016b), see Table 2 in Appendix B. We segment the dataset into non-overlapping 1,024 token windows (the model context limit) and compute SBERT embedding for every window. We cluster the generalist dataset hierarchically with a clustering tree with branching $64$ for $4$ levels, see Appendix A. The levels therefore have 64, 4,096 ($=64^{2}$), 260k ($=64^{3}$) and 16.7m ($=64^{4}$) clusters with an average of 540B, 8.4B, 130m and 2m tokens per cluster respectively. As an alternative to SBERT embeddings, we also consider Latent Semantic Index (LSI), i.e. singular value decomposition over tf-idf representations (Deerwester et al., 1990; Dumais, 2004).
For our specialist tasks, we consider 3 language modeling tasks (LM) and 3 multiple-choice-question tasks (MCQ). For LM, we use Pile subsets from different domains (Gao et al., 2021): medical (Pubmed Central), programming Q&A (Stackexchange), and encyclopedic (Wikipedia). For MCQ answering, we use AI2 Reasoning Challenge (Clark et al., 2018, ARC), Massive Multitask Language Understanding (Hendrycks et al., 2021, MMLU), and Reward Bench Reasoning (Lambert et al., 2024, RWDB-R). ARC focuses on science questions, MMLU focuses on interdisciplinary knowledge, RWDB-R focuses on correct vs incorrect solutions to math and programming problems. To provide a representative specialist train set $D^{\rm s}\sim{\cal D}^{\rm s}$, we split the questions into a train and test split, see Table 3 in Appendix B.
Our main results are reported with unified settings. For the classifier, the classification threshold is the main parameter. A threshold accepting 2.5% of $D^{\rm g}$ worked best in for the runs with 1.3B models over 120B tokens. For DoGE, the method is costly to apply over many data sources/clusters and we applied it over 64 clusters, i.e. learning a mixture weight of dimension 64. For importance sampling, the results presented in this section relies on 260k clusters. Later, Section 5 studies ablations and parameter sensitivity. Details on hyperparameters can be found in Appendix C.
@@ 4.1 Language Modeling Tasks
We evaluate specialist LMs on three domains from the Pile (Gao et al., 2021): medical (PubMed Central), encyclopedic (Wikipedia) and programming Q&A (StackExchange). We limit specialist training data from 14m tokens to the full Pile subset, up to 26.7B tokens, see Table 2 in Appendix B.
@@
We report pretraining results in Figure 4, and the fine-tuning results in Figure 4. For each domain, the pretraining results evaluate models trained using the resampled Redpj2 examples. The fine tuning results evaluate models where each model pretrained on (resampled) Redpj2 has been further trained on the in-domain data itself (PubMed, StackExchange, Wikipedia). All experiments consider the same optimization effort and we validate the fraction of steps spent in fine-tuning, from 3%-ft with 14m tokens to 100%-ft with 26.7B tokens.
The pretraining results in Figure 4 show that the in-domain perplexity is better with task-dependent pretraining than with generic pretraining (base Redpj2) for all methods. This gain in perplexity comes as model training focuses on data close to the targeted domain: the model capacity is not used to fit the filtered out training data. Table 8 in Appendix E shows, for instance, that CRISP outperforms base on 97.3% of PubMed but reports worse perplexity on 95.9% of Redpj2.
When we fine tune the pretrained models, the advantage of task-dependent pretraining is preserved, in Figure 4. Task-specific pretraining checkpoints are better starting points for fine-tuning than generic ones. This shows the complementarity between task-dependent pretraining and fine-tuning. Figure 4 also shows the necessity of pretraining: below 7B tokens, the “only specific” 1.3B model shows high perplexity. When comparing task-dependent pretraining methods, importance sampling consistently performs better after fine-tuning, even when the pretraining results are close (e.g. classifier on PubMed, Wikipedia).
@@ 4.2 Multiple Choice Question Tasks
Compared to LM, MCQ has much smaller specialist training sets per task, i.e. between 200k and 2m tokens, see Table 3 in Appendix B. The MCQ evaluation is also different: it uses accuracy and not perplexity. For each example, the model scores the concatenation of the question and a possible answer, for each proposed answer. The model is accurate when the correct answer is assigned the highest score (probability or normalized probability, see Appendix D). For MCQ tasks, unlike for LM tasks, the training loss (negative log likelihood) is therefore not closely tied to the test metric.
Despite these differences, we observe a similar benefit for task-dependent pretraining compared to task-agnostic (base) pretraining. Figure 4 displays a similar method ordering and CRISP is consistently the best method. As a difference with LM tasks, we observe limited benefits from fine tuning, see Figure 4. Fine-tuning improves the base method on all datasets except ARC-E, but not enough to outperform task-specific pretraining, see Table 10 in Appendix F.

### v2

# Task-Adaptive Pretrained Language Models via Clustered Importance Sampling

@@ Abstract
Specialist language models (LMs) focus on a specific task or domain on which they often outperform generalist LMs of the same size. However, the specialist data needed to pretrain these models is only available in limited amount for most tasks. In this work, we build specialist models from large generalist training sets instead. We propose a novel method, ClusteRed Importance SamPling (CRISP). CRISP clusters the generalist dataset and samples from these clusters based on their frequencies in the smaller specialist dataset. It is scalable, suitable for both pretraining and continued pretraining, and works well in multi-task settings. CRISP performs favorably compared to other methods that adjust the training distribution of the generalist data with guidance from the limited domain-specific data. Our findings demonstrate improvements across different domains in terms of language modeling perplexity and accuracy on multiple-choice question tasks. We also present ablation studies that examine the impact of dataset sizes, clustering configurations, and model sizes.
@@ 1 Introduction
This paper presents an exhaustive comparison over different model sizes (350m, 1.3B, 6.8B) and different numbers of clusters (scaling from 64 to 16m clusters with hierarchical clustering). We consider different tasks, both for language modeling and multiple-choice questions. We also explain the impact of hyperparameters such as the clustering representation and number of clusters. We study IS in the context of multitasking and continued pretraining. We also perform ablations with respect to the generic pre-training set size and the specialization data size.
@@ 2 Related Work
Generalist vs Specialist LMs Generalist LMs address tasks for which they have not been explicitly trained (Brown et al., 2020) or provide a good initialization for fine-tuning a dedicated model (Devlin et al., 2019). Nowadays generalists compete with dedicated models on many tasks (Jiang et al., 2024; Dubey et al., 2024). Success, however, comes at a price: a generalist must be much larger than a specialist for the same accuracy. For instance, on English-to-German translation, the 175-B parameter generalist GPT-3 (Brown et al., 2020) is less accurate than a 136m-parameter specialist (Sennrich et al., 2016a). For neural LMs, the parameter count directly impacts training and inference costs.
Specialist large LMs exist in domains where large amounts of specialized texts are available. Corpora with billions of tokens enable pretraining or continued pretraining, a generalist pretraining phase followed by a specialist one (Gururangan et al., 2020; Parmar et al., 2024)). Domains with specialist models include medicine and biology (Lewis et al., 2020; Labrak et al., 2024; Bolton et al., 2024), computer programming and mathematics (Lewkowycz et al., 2022; Rozière et al., 2024; Azerbayev et al., 2024) and finance (Wu et al., 2023; Xie et al., 2023a). When specialist data is available in limited amount, task-adaptive data-selection methods train specialist models on generalist data instead.
Task-Adaptive Data-Selection These selection methods over-sample generalist data that aids model generalization in the target domain. For masked LMs, Gururangan et al. (2020) observe that continued pretraining improves the performance on end-tasks when using data with high vocabulary overlap with the targeted task. For machine translation (MT), Aharoni & Goldberg (2020) show that a task-adapted pretraining dataset can be selected from a generalist dataset using the nearest neighbors of a small specialist set. Their nearest neighbor classifier relies on BERT sentence distance (Devlin et al., 2019). Still for MT, other works have used other types of classifiers. In particular, contrasting the scores of two LMs (generalist and specialist) is popular (Moore & Lewis, 2010; Axelrod et al., 2011; Wang et al., 2018; Junczys-Dowmunt, 2018). Other classifiers include logistic regression or fine-tuned BERT (Iter & Grangier, 2021). Outside classification, Xie et al. (2023c) proposed to use importance sampling for continued pretraining. They improve classification tasks by selecting pretraining data with a similar distribution to the targeted domain in terms of hashed-ngrams. Importance sampling is also used in (Grangier et al., 2024b) and we build upon that work which adjusts the frequency of generalist clusters informed by specialist data: we scale the method to millions of clusters, show that it works with larger models, and extend it beyond language modeling tasks.
@@ 3.1 Notations
$D^{\rm g}$ is the training dataset sampled from the generalist distribution ${\cal D}^{\rm g}$. $D^{\rm s}$ is the specialist dataset representative of the final task, sampled from the specialist distribution ${\cal D}^{\rm s}\neq{\cal D}^{\rm g}$. The loss of model $\theta$ on a dataset $D$ is
@@
where $|D|$ denotes the cardinality of $D$ and $|x|$ denotes the length of sequence $x=(x_{1},\ldots,x_{|x|})$. The perplexity of model $\theta$ on the dataset $D$ is ${\cal P}(D;\theta):=\exp({\cal L}(D;\theta)).$
@@ 3.2 Classification
which restricts the generic dataset $D^{\rm g}$ to the examples with an estimated probability to be in-domain above threshold $t$. The threshold $t$ is a sensitive hyperparameter that impacts the downstream model. It is validated as a trade-off between focusing on data close to the domain of interest while keeping $C(D^{\rm g},t)$ large enough to train an LM of the targeted capacity. In our case, we rely on a logistic regression classifier trained over sentence BERT (SBERT) text embeddings (Reimers & Gurevych, 2019), an established classification method (Minaee et al., 2021). The SBERT representation is also commonly used in data selection (Albalak et al., 2024; Xie et al., 2023c; Zhang et al., 2024; Su et al., 2023). This representation is also used in the alternative selection strategies we consider. As an ablation, we also evaluate the impact of the choice of SBERT (Section 5.1).
@@ 3.4 CRISP: ClusteRed Importance Sampling for Pretraining
Figure 1: Task-adaptive data selection with Clustered Importance Sampling (CRISP).
@@
${\cal L}({\cal D}^{\rm s};\theta)=\sum_{c}{\cal L}(c;\theta)P(c|{\cal D}^{\rm s})=\sum_{c}{\cal L}(c;\theta)\frac{P(c|{\cal D}^{\rm s})}{P(c|{\cal D}^{\rm g})}P(c|{\cal D}^{\rm g})=\mathop{\mathbb{E}}_{c\sim(c|{\cal D}^{\rm g})}[w(c){\cal L}(c;\theta)].$
In our experiments, we estimate the terms $w(c),{\cal L}(c;\theta)$ from the finite training sets $D^{\rm s}\sim{\cal D}^{\rm s}$ and $D^{\rm g}\sim{\cal D}^{\rm g}$. We count the number of examples in each cluster to estimate $P(c|{D}^{\rm s}),P(c|{D}^{\rm g})$. The expected loss over a cluster ${\cal L}(c;\theta)$ is estimated as the average loss over the generalist examples in cluster $c$, ${\cal L}({D}^{\rm g}\cap K(c);\theta)$, where $K(c)$ denotes the examples in cluster $c$. This strategy therefore only estimates $P(c|{D}^{\rm s})$ on the small ${D}^{\rm s}$. The term ${\cal L}(c;\theta)$ is estimated over the large set as ${\cal L}({D}^{\rm g}\cap K(c);\theta)$ which uses many more samples and hence has less variance than the estimator ${\cal L}({D}^{\rm s}\cap K(c);\theta)$ over the small ${D}^{\rm s}$.
We train CRISP models with stochastic optimization (Kingma & Ba, 2015, Adam) and propose Algorithm 1. Here, we do not explicitly reweigh the loss. We instead sample clusters from their importance. This avoids frequently visiting clusters with less weight. This strategy has less variance in its gradient estimates, which can help convergence (Seiffert et al., 2008; An et al., 2021). This algorithm is simple and efficient when one groups the generalist examples by cluster prior to training.
@@
8:         $x_{i}\sim{\rm Uniform}({D}^{\rm g}\cap K(c))$triangleright Sample a generalist example in the selected cluster.
@@
10:     $\theta_{t}\leftarrow{\rm AdamUpdate}(\theta_{t-1},\{x_{1},\ldots,x_{B}\})$
@@ 4 Experiments & Results
We perform experiments with transformer LMs (Vaswani et al., 2017). Most of our experiments use models with 1.3B parameters (trained on 120B tokens) and we conduct ablations with 350m and 7B models (resp. trained on 40B, 350B tokens). Our settings for architectures and optimization are borrowed from Brown et al. (2020), see Appendix D.
Our generalist training set is Redpj2 (Together AI Team, 2023). We select this dataset as it contains only web-crawled data without additional interventions to help evaluation tasks (e.g. adding encyclopedias, books or academic articles). Redpj2 contains over 30T tokens with our 32k byte-pair encoding tokenizer (Sennrich et al., 2016b), see Table 4 in Appendix C. We segment the dataset into non-overlapping 1,024 token windows (the model context limit) and compute SBERT embedding for every window. We cluster the generalist dataset hierarchically with a clustering tree with branching $64$ for $4$ levels, see Appendix B. The levels therefore have 64, 4,096 ($=64^{2}$), 260k ($=64^{3}$) and 16.7m ($=64^{4}$) clusters with an average of 540B, 8.4B, 130m and 2m tokens per cluster respectively. As an alternative to SBERT embeddings, we also consider Latent Semantic Index (LSI), i.e. singular value decomposition over tf-idf representations (Deerwester et al., 1990; Dumais, 2004).
For our specialist tasks, we consider 3 language modeling tasks (LM) and 3 multiple-choice-question tasks (MCQ). For LM, we use Pile subsets from different domains (Gao et al., 2021): medical (Pubmed Central), programming Q&A (Stackexchange), and encyclopedic (Wikipedia). For MCQ answering, we use AI2 Reasoning Challenge (Clark et al., 2018, ARC), Massive Multitask Language Understanding (Hendrycks et al., 2021, MMLU), and Reward Bench Reasoning (Lambert et al., 2024, RWDB-R). ARC focuses on science questions, MMLU focuses on interdisciplinary knowledge, RWDB-R focuses on correct vs incorrect solutions to math and programming problems. To provide a representative specialist train set $D^{\rm s}\sim{\cal D}^{\rm s}$, we split the questions into a train and test split, see Table 5 in Appendix C.
Our main results are reported with unified settings. For the classifier, the classification threshold is the main parameter. A threshold accepting 2.5% of $D^{\rm g}$ worked best in for the runs with 1.3B models over 120B tokens. For DoGE, the method is costly to apply over many data sources/clusters and we applied it over 64 clusters, i.e. learning a mixture weight of dimension 64. For importance sampling, the results presented in this section relies on 260k clusters. Later, Section 5 studies ablations and parameter sensitivity. Details on hyperparameters can be found in Appendix D.
@@ 4.1 Language Modeling Tasks
We evaluate specialist LMs on three domains from the Pile (Gao et al., 2021): medical (PubMed Central), encyclopedic (Wikipedia) and programming Q&A (StackExchange). We limit specialist training data from 14m tokens to the full Pile subset, up to 26.7B tokens, see Table 4 in Appendix C.
@@
We report pretraining results in Figure 4, and the fine-tuning results in Figure 4. For each domain, the pretraining results evaluate models trained using the resampled Redpj2 examples. The fine tuning results evaluate models where each model pretrained on (resampled) Redpj2 has been further trained on the in-domain data itself (PubMed, StackExchange, Wikipedia). All experiments consider the same optimization effort and we validate the fraction of steps spent in fine-tuning, from 3%-ft with 14m tokens (97% pretraining) to 100%-ft with 26.7B tokens (no pretraining).
The pretraining results in Figure 4 show that the in-domain perplexity is better with task-dependent pretraining than with generic pretraining (base Redpj2) for all methods. This gain in perplexity comes as model training focuses on data close to the targeted domain: the model capacity is not used to fit the filtered out training data. Table 10 in Appendix F shows, for instance, that CRISP outperforms base on 97.3% of PubMed but reports worse perplexity on 95.9% of Redpj2.
When we fine tune the pretrained models, the advantage of task-dependent pretraining is preserved, as shown in Figure 4. Task-specific pretraining checkpoints are better starting points for fine-tuning than generic ones. This shows the complementarity between task-dependent pretraining and fine-tuning. Figure 4 also shows the necessity of pretraining: below 7B tokens, the “only specific” 1.3B model shows high perplexity. When comparing task-dependent pretraining methods, importance sampling consistently performs better after fine-tuning, even when the pretraining results are close (e.g. classifier on PubMed, Wikipedia).
@@ 4.2 Multiple Choice Question Tasks
Compared to LM, MCQ has much smaller specialist training sets per task, i.e. between 200k and 2m tokens, see Table 5 in Appendix C. The MCQ evaluation is also different: it uses accuracy and not perplexity. For each example, the model scores the concatenation of the question and a possible answer, for each proposed answer. The model is accurate when the correct answer is assigned the highest score (probability or normalized probability, see Appendix E). For MCQ tasks, unlike for LM tasks, the training loss (negative log likelihood) is therefore not closely tied to the test metric.
Despite these differences, we observe a similar benefit for task-dependent pretraining compared to task-agnostic (base) pretraining. Figure 4 displays a similar method ordering and CRISP is consistently the best method. As a difference with LM tasks, we observe limited benefits from fine tuning, see Figure 4. Fine-tuning improves the base method on all datasets except ARC-E, but not enough to outperform task-specific pretraining, see Table 12 in Appendix G.